OpenStudy (anonymous):

Simplify. Please show all of the steps to arrive at your answer. 33 ÷ 9 ∙ 2 – 12 + 6

razor99 (razor99):

waht is that dot bro

OpenStudy (anonymous):

multiply sign

razor99 (razor99):

first do u know bodmas rule

razor99 (razor99):

helloo u there

OpenStudy (anonymous):

no i dont

razor99 (razor99):

okay Bodmas rule is like this B=bracket o=of d=division m=multiplication a=addition s=subtraction Now looking a the question the rules say first u have to slove the brackets their is no bracket so we have to look for the division sign.

razor99 (razor99):

i am typing the answer right now

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im still lost

OpenStudy (anonymous):

1.33?

razor99 (razor99):

wait thats no the answer its the first step

OpenStudy (anonymous):

oh and it is suppose to be 3 to the third power not 33

razor99 (razor99):

hey u confused me its 3 right

razor99 (razor99):

3÷9x2-12+6 =3x2-12+6 =6-12+6 =6-18 =-12

OpenStudy (anonymous):

simple PEMDAS please